OpenStudy (anonymous):

determine whether the series converges absolutely, converges conditionally or diverges

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ (-1)^n *(2n)!}{ 5n! }\]

OpenStudy (anonymous):

i got an = (2n)!/5n! i dont know what the ! symbols mean

OpenStudy (anonymous):

means "factorial"

OpenStudy (anonymous):

what do i do with it?

OpenStudy (anonymous):

like how do i determine if converges absolutely or conditionally?

OpenStudy (anonymous):

nothing really \((2n)!\) is much larger than \(5n!\) the terms do not even go to zero

OpenStudy (xapproachesinfinity):

this is a divergent series take the limit

OpenStudy (anonymous):

so if they dont go to zero then that means that it diverges?

OpenStudy (anonymous):

@satellite73 like this \[\lim_{n \rightarrow \infty} \frac{ (2n)! }{ 5n! }\]

OpenStudy (anonymous):

@rational could help me with this?

OpenStudy (anonymous):

what do i with the factorial symbol?

OpenStudy (rational):

are you sure the denominator is 5*n! and not (5n)! ?

OpenStudy (anonymous):

yeah its 5n!

OpenStudy (rational):

take a screenshot and attach maybe..

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

here it is @rational

OpenStudy (rational):

yes it does not converge you may use ratio test to conclude that

OpenStudy (anonymous):

we havent covered ratio test yet :(

OpenStudy (anonymous):

for this other one \[\frac{ 16^\frac{ 1}{ 2 } }{ 7^{n+1}}\]

OpenStudy (anonymous):

what do i do with the n+1 exponent?

OpenStudy (rational):

you want to know whether it converges or diverges ?

OpenStudy (anonymous):

here is the whole problem

OpenStudy (anonymous):

no whether it coverges absolutely conditionally or diverges?

OpenStudy (rational):

simplify this first \[\frac{16^{n/2}}{7^{n+1}}\]

OpenStudy (rational):

\[\frac{16^{n/2}}{7^{n+1}} = \frac{(4^2)^{n/2}}{7\cdot 7^{n}} = \dfrac{1}{7}\cdot \frac{4^n}{7^n} = \frac{1}{7}\cdot \left(\frac{4}{7}\right)^n\]

OpenStudy (rational):

does that look familair ?

OpenStudy (anonymous):

sort of looks like a harmonic series right?

OpenStudy (rational):

\[\sum\limits_{n=1}^{\infty} (-1)^{n-1}\frac{16^{n/2}}{7^{n+1}} ~~=~~\frac{1}{7}\sum\limits_{n=1}^{\infty} (-1)^{n-1}\left(\frac{4}{7}\right)^{n} \]

OpenStudy (rational):

it is a geometric series with a common ratio of \(\frac{4}{7}\)

OpenStudy (rational):

does the geometric series with common ratio 4/7 converges or diverges ?

OpenStudy (anonymous):

converges right?

OpenStudy (rational):

Yep! and it converges absolutely

OpenStudy (anonymous):

how do you determine its absolutely?

OpenStudy (rational):

because the absolute value of the series : \[\left|\sum\limits_{n=1}^{\infty} (-1)^{n-1}\frac{16^{n/2}}{7^{n+1}} \right|\] is same as geometric series which converges

OpenStudy (rational):

if the absolute value of a series converges, then the original series also converges. then we say that the "series converges absolutely"

OpenStudy (anonymous):

and to see if the absolute value of the series converges we find its limit right?

OpenStudy (rational):

what we found above is that absolute value of the given series converges

OpenStudy (rational):

\[\left|\sum\limits_{n=1}^{\infty} (-1)^{n-1}\frac{16^{n/2}}{7^{n+1}} \right|~~~ =~~~\sum\limits_{n=1}^{\infty}\frac{16^{n/2}}{7^{n+1}} \]

OpenStudy (rational):

that right side is a converging geometric series because the ratio is 4/7 that proves the absolute value of given series converges

OpenStudy (anonymous):

oh got it

OpenStudy (anonymous):

for the last problem with the factorial symbols how did we proof it diverged again?

OpenStudy (anonymous):

thanks for your help @rational